OpenStudy (ellecullen):

Homework Question: solve the following equation: sin x cos x = sqrt3/4

OpenStudy (mathmale):

The first way in which i'd try solving this problem would be to re-write the left side. Look for a trig identity that involves sin x cos x.

OpenStudy (ellecullen):

I did this: sqrt3/4 = 1/2 x sqrt3/2 sqrt3/2= cos(pi/6) then sin(x) x sqrt3/2) = sqrt3/4 sin(x) = (sqrt3/4) / (sqrt3/2) sin(x) = 1/2

OpenStudy (mathmale):

Have you found the value of x? In either case, how many solutions could you expect? 1? 2? 3? Have you checked your answer(s)?

OpenStudy (mathmale):

I do agree that your sin x = 1/2 leads to a solution. But what is that solution? And, again, is this the only solution?

OpenStudy (ellecullen):

My solution is pi/6 according to the unit circle stuff. It looks as if there is one solution: pi/6

OpenStudy (mathmale):

Would you mind checking whether or not 5Pi/6 is a solution?

OpenStudy (ellecullen):

5pi/6 is a solution for sin(x) = 1/2 meanwhile its cosine is a -(sqrt3/2). In this case, the -(sqrt3/2) does not work.

OpenStudy (mathmale):

So, I hear you saying that Pi/6 is a solution and 5Pi/6 is not. Is that what you meant?

OpenStudy (ellecullen):

yes

myininaya (myininaya):

Also sin(x)cos(x) can be written as 1/2 *2sin(x)cos(x)=1/2 * sin(2x)

hero (hero):

\[\sin(x)\cos(x)=\frac{\sqrt{3}}{4}\] \[4\sin(x)\cos(x)=\sqrt{3}\] \[(4\sin(x)\cos(x))^2=(\sqrt{3})^2\] \[16\sin^2(x)\cos^2(x)=3\] \[16\sin^2(x)(1−\sin^2(x))=3\] \[16\sin^2(x)−16\sin^4(x)=3\] Let \(y = \sin^2(x)\) \(16y−16y^2=3\) \(−16y^2+16y−3=0\) \(16y^2−16y+3=0\) \(16y^2−(12+4)y+3=0\) \(16y^2−12y−4y+3=0\) \(4y(4y−3)−1(4y−3)=0\) \((4y−3)(4y−1)=0\) \(4y−3=0\) \(4y−1=0\) Replace y with \(\sin^2(x)\) \(4\sin^2(x)−3=0\) \[4\sin^2(x)−1=0\]

hero (hero):

\[\sin^2(x) = \frac{3}{4}\] \[\sin^2(x) = \frac{1}{4}\] \[\sin(x) = \pm \frac{\sqrt{3}}{4}\] \[\sin(x) = \pm \frac{1}{2}\] only positive solutions for x